OpenStudy (anonymous):

Use the equation mg sinθ = u mg cos θ to determine the steepest slope of a street on which a car with rubber tires can park without sliding.

OpenStudy (anonymous):

\[\mu=\frac{ \sin \theta }{ \cos \theta }=\tan \theta\]

OpenStudy (anonymous):

So, \[\theta=\tan^{-1} \mu\]

OpenStudy (anonymous):

The mg on each side cancels out